Abstract

Elena Amanda Long, Ph.D., August, 2012 Physics POLARIZED ^{3}He(e, e′n) ASYMMETRIES IN THREE ORTHOGONAL MEASUREMENTS Directors of Dissertation: Bryon D. Anderson and Douglas W. Higinbotham

Asymmetry measurements were conducted in Jefferson Lab’s experimental Hall A through electron scattering from a polarized^{3}He target in the quasi-elastic^{3}He(e, e′n) reaction. Measurements were made with the target polarized in the longitudinal direction with respect to the incoming electrons (A_{L}), in a transverse direction that was orthogonal to the beam-line and parallel to the q-vector (A_{T}), and in a vertical direction that was orthogonal to both the beam-line and the q-vector (A_{y}^{0}). The experiment measured A_{y}^{0} at four-momentum transfer squared (Q^{2}) of 0.127 (GeV/c)^{2}, 0.456 (GeV/c)^{2} and 0.953 (GeV/c)^{2}. The A_{T} and A_{L} asymmetries were both measured at Q^{2} of 0.505 (GeV/c)^{2} and 0.953 (GeV/c)^{2}. This is the first time that three orthogonal asymmetries have been measured simultaneously. Results from this experiment are compared with the plane wave impulse approximation (PWIA) and Faddeev calculations. These results provide important tests of models used to extract neutron form factors from polarized ^{3}He.